OpenStudy (anonymous):

Either prove or provide a counterexample to the assertion "If n=ab is the product of positive integers a and b then a <= the root of n or b <= the root of n."

OpenStudy (anonymous):

n = 3 * 3 n = 9 √9 = 3 so this proves it's true but n = 100 * 1 n = 100 √100 = 10 so this situation proves that it is false. Once you get a false statement the entire thing is false so there.